|5.9-4.7x|+2.1=17.4
Answer by lwsshak3(11628)

|5.9-4.7x|+2.1=17.4
solve for two conditions:
for (5.9-4.7x)>0
5.9-4.7x+2.1=17.4
-4.7x=9.4
x=-9.4/4.7=2
..
for (5.9-4.7x)<0
-5.9+4.7x+2.1=17.4
4.7x=21.2
x=21.2/4.7≈4.51